OpenStudy (anonymous):

Find all solutions to the equation. 2cos(theta) + sqrt 2 = 0

OpenStudy (ash2326):

let's represent cos theta by x, can you find value of x? \[2x+\sqrt 2=0\]

OpenStudy (anonymous):

-√2/2

OpenStudy (ash2326):

Good or it can be simplified as \[\frac{-1}{\sqrt 2}\]

OpenStudy (anonymous):

oh ok

OpenStudy (ash2326):

We have now \[\cos \theta=\frac{-1}{\sqrt 2}\]

OpenStudy (ash2326):

do you know the value of theta if cos theta =1 /sqrt 2

OpenStudy (anonymous):

mmm no

OpenStudy (ash2326):

what's cos 45?

OpenStudy (anonymous):

0.52532198881 ?

OpenStudy (anonymous):

i just plugged that in my calculator :/

OpenStudy (anonymous):

Hmmm, there might be another way than just calculator.

OpenStudy (anonymous):

oh i see.

OpenStudy (anonymous):

how to I write the answer in radians or in terms of pie?

OpenStudy (anonymous):



OpenStudy (anonymous):

It's gotta be half or a right angle to work.

OpenStudy (anonymous):

A right angle in radians is \(2\pi / 4 = \pi / 2\). So what is half of that?

OpenStudy (anonymous):

There are 4 right angles in a full revolution.

OpenStudy (anonymous):

half of π/2? i don't know